OpenStudy (superhelp101):

What are the vertices of the ellipse given by the equation 4x2 + y2 - 80x - 6y + 309 = 0? (-20, 3) and (0, 3) (10, -13) and (10, 7) (20, 3) and (0, 3) (10, 13) and (10, -7)

OpenStudy (superhelp101):

@study100

OpenStudy (superhelp101):

I think it's D ?

OpenStudy (scorcher219396):

You'll need to put this into graphing form by completing the square \[4x^2+y^2-80x-6y+309=0\] Start by rearranging so the x's are together and the y's are together \[4x^2-80x+y^2-6y=-309\] Factor out the coefficients \[4(x^2-20x+Q)+(y^2-6y+P)=-309\] I left Q and P in there because you need to then add constants to both sides that will make what is in the parentheses a perfect square 4(x^2-20x+100)+(y^2-6y+9)=-309+400+9 4(x-10)^2+(y-3)^2=100 Divide everything by 100 to get a 1 on the right \[\frac{ (x-10)^2 }{ 25 }+\frac{ (y-3)^2 }{ 100 }\] So your center is at (10,3), the vertices at (5,-3), (15,-3), (10,-13), and (10,7) So B, not D

OpenStudy (anonymous):

yes you're correct! :)

OpenStudy (superhelp101):

awww thanks! you're very sweet for writing it out! :)

OpenStudy (superhelp101):

I am going to cry right now. You're so kind :)

OpenStudy (anonymous):

hm? k and a are positive @scorcher219396

OpenStudy (superhelp101):

that's is what I noticed too..:)

OpenStudy (anonymous):

well glad it helps :) ... ily.. too! :))

OpenStudy (superhelp101):

hehehehe :))))))))) <3